Calculate the heat change in kilocalories for vaporization of 49 g of water at 100 C.

Please express answer as a positive value using two significant figures and include the appropriate units.
1
Expert's answer
2020-12-18T08:36:34-0500

The heat of vaporization at "100 \ ^o C = 540\ \frac{cal}{g} = 0.54\ \frac{kcal}{g}"

The absorbed heat "= 49 \ g * 0.54\ \frac{kcal}{g} = 26.46\ kcal \approx 26\ kcal"


Answer: "26\ kcal"